I'm really confused with the process that I'm supposed to follow. I am finding it difficult to even know where to start.
The first problem is:
-x+y=2
_______
x+y=4
The only thing I know for sure is that I'm supposed to use the substitution method but what am I substituting, with what and why.
I really do appreciate whatever help you can give me.
Thank you

Answer by stanbon(75887)

-x+y=2
_______
x+y=4
The only thing I know for sure is that I'm supposed to use the substitution method but what am I substituting, with what and why.
-------
It's like saying "my brother is 2 years older than I am".
If the person knows how old you are they can figure out
how old your brother is.
---------
So you solve one of the equations for one of the variables:
If x+y=4 then x=4-y
Now we know how y is related to x.
So, inplace of x in the other equation we put what x is equal to, as follows:
-(4-y)+y=2
Then we can figure out what Y is, as follows:
-4+y+y=2
-4+2y=2
2y=6
y=3
Now we know what y equals, so substitute back and find out what x is:
If x=4-y then x=4-3=1
Now we know that x=1 and y=3
Tha's it; you've found the value of both variables for these two equations:
Cheers,
Stan H.